Let A be the event that a student is enrolled in an accounting course, and let S be the event that a student is enrolled in a statistics course. It is known that 30% of all students are enrolled in an accounting course and 40% of all students are enrolled in statistics. Included in these numbers are 15% who are enrolled in both statistics and accounting. A student is randomly selected, what is the probability that the student is enrolled in either accounting or statistics or both?
Answer by Boreal(15235)

P(S OR A)=P(S)+P(A)-P(S AND A); the last takes care of double counting.
=0.3+0.4-0.15=0.55.